A rectangle with the dimensions of 2 feet by 8 feet is similar to a rectangle with the dimensions of
Answer by jim_thompson5910(35256)

It's similar to any rectangle with dimensions of W and L such that L = 4W (length is 4 times the width). This is because 8/2 = 4, so the ratio of the two dimensions is 1:4 (W:L)


So if W = 3, then L = 4*W = 4*3 = 12. One rectangle is 3 by 12

Or, if W = 5, then L = 4*W = 4*5 = 20. Another rectangle is 5 by 20.

There are infinitely many rectangles.